# Mean Field Game of Optimal Tracking Portfolio

### v2

@@ 3 Equivalent Problem Formulation for Best Response Control
To tackle the best response control of the representative agent in problem (2.5), we first reformulate the problem based on the observation that, for a fixed control $\theta$, the optimal capital injection $A_{t}^{\theta,*}$ is always the smallest adapted right-continuous and non-decreasing process that dominates $\lambda\overline{V}_{t}+(1-\lambda)Z_{t}-V_{t}$ for $t\in[0,T]$. It follows from Lemma 2.4 in Bo et al. (2021) that, for a fixed regular control $\theta=(\theta_{t})_{t\in[0,T]}$, the optimal singular control satisfies $A^{\theta,*}_{t}=0\vee\sup_{s\leq t}\left(\lambda\overline{V}_{s}+(1-\lambda)Z_{s}-V_{s}\right)$, $\forall t\in[0,T]$. In particular, the initial capital injection $A^{\theta,*}_{0}=(1-\lambda)(z-\mathrm{v})^{+}$ with $(x)^{+}:=\max\{x,0\}$ for $x\in\mathds{R}$. Thus, the control problem (2.5) admits the equivalent formulation as an unconstrained control problem involved a running maximum cost that
${\rm w}(\mathrm{v},z)=\inf_{\theta\in\mathbb{U}^{\rm r}}\mathbb{E}\left[A_{0}^{\theta,*}+\int_{0}^{T}e^{-\rho t}dA_{t}^{\theta,*}\right]$
$=(1-\lambda)(z-\mathrm{v})^{+}+\inf_{\theta\in\mathbb{U}^{\rm r}}\ \mathbb{E}\left[\int_{0}^{T}e^{-\rho t}d\left(0\vee\sup_{s\leq t}\left(\lambda\overline{V}_{s}+(1-\lambda)Z_{s}-V_{s}\right)\right)\right],$ (3.1)
where $\mathbb{U}^{\rm r}$ is the admissible control set of portfolio strategies $\theta=(\theta_{t})_{t\in[0,T]}$ such that $\theta$ is an $\mathbb{F}$-adapted real-valued process satisfying $\mathbb{E}[\int_{0}^{T}|\theta_{t}|^{2}dt]<\infty$.
Remark 3.1.
We note that the capital injection by the fund manager in reality may not be possible or allowed. This is the reason that we call the singular control $A=(A_{t})_{t\in[0,T]}$ a “fictitious capital injection” in the present paper, because it can be interpreted as a capital injection or bail-out strategy such that the total wealth $V_{t}+A_{t}$ stays above the benchmark process for $t\in[0,T]$, which is similar to the role of actual capital injection in the literature of dividend payment, for instance, see Eisenberg and Schmidli (2009) and Lokka and Zervos (2008). But our problem can also be equivalently formulated as a stochastic control problem under the running maximum cost. As we have $A_{t}^{\theta,*}=0\vee\sup_{s\leq t}\left(\lambda\overline{V}_{s}+(1-\lambda)Z_{s}-V_{s}\right)$, the term $\mathbb{E}[\int_{0}^{T}e^{-\rho t}dA_{t}^{\theta,*}]$ can also be understood as a risk measure on the expected largest shortfall with reference to the benchmark $\lambda\overline{V}_{s}+(1-\lambda)Z_{s}$, see also the conventional definition of expected shortfall with respect to a random variable at the terminal time in Pham (2002) and the intra-horizon expected shortfall considered in Farkas et al. (2021) and references therein.
To solve the stochastic control problem (3), we introduce a new controlled state process to replace the original state process. Let $K_{t}:=\lambda\overline{V}_{t}+(1-\lambda)Z_{t}-V_{t}$ and $L_{t}:=0\vee\sup_{s\leq t}K_{s}$ for $t\in[0,T]$. Then, we define a new state process with reflection that $X_{t}:=L_{t}-K_{t}$, and hence
@@
$u(t,x,z):=\sup_{\theta\in\mathbb{U}_{t}^{r}}\mathbb{E}\left[-\int_{t}^{T}e^{-\rho(s-t)}dL_{s}^{X}\big|X_{t}=x,Z_{t}=z\right],$ (3.3)
where the admissible control set $\mathbb{U}_{t}^{\rm r}$ is specified as the set of $\mathbb{F}$-adapted real-valued control processes $\theta=(\theta_{s})_{s\in[t,T]}$ satisfying $\mathbb{E}[\int_{t}^{T}|\theta_{s}|^{2}ds]<\infty$. Obviously, we have $\mathbb{U}^{\rm r}=\mathbb{U}^{\rm r}_{0}$ and the following equivalence between problems (3) and (3.3):
@@
We refer to $\mathcal{O}_{T}$ as the capital injection region and $\mathcal{O}_{T}^{c}\cap\overline{\mathcal{D}}_{T}$ as the non-capital injection region. In the capital injection region, even under the optimal portfolio strategy, the agent’s wealth requires external capital injections to meet or exceed the benchmark process. In contrast, in the non-capital injection region, the agent can adopt a portfolio strategy such that the resulting wealth process can always outperform the benchmark without any capital injections. Moreover, Theorem 3.4 yields that the next corollary on the semi-explicit form of the optimal portfolio (feedback) function given by
@@ 4 Mean Field Equilibrium
and establishes the existence of a positive function $f^{*}\in{\tt{C}}([0,T])$ satisfying (4.1). Consequently, the MFE is given by $(f^{*},\theta^{*,f^{*}},A^{*,f^{*}})=(f^{*}(t),\theta_{t}^{*,f^{*}},A_{t}^{*,f^{*}})_{t\in[0,T]}$ characterized in Definition 2.1. On the other hand, it follows from (2.6) that $V_{t}^{\theta^{*,f}}=\mathrm{v}+\int_{0}^{t}\theta_{s}^{*,f}\mu ds+\int_{0}^{t}\theta_{s}^{*,f}\sigma dW_{s}$ for $t\in[0,T]$. Then, taking the expectation on both sides of this equation, we have
@@
Lemma 4.2 shows that the capital injection region $\cal O$} and non-capital injection region ${\cal O}_{T}^{c}\cap\overline{\mathcal{D}}_{T}$ are not interconnected. In other words, if the initial state $(0,x,z)\in{\cal O}_{T}$, $(t,X_{t}^{*},Z_{t})\in{\cal O}_{T}$, $\mathbb{P}$-a.s., for any $t\in[0,T]$; if $(0,x,z)\in{\cal O}_{T}^{c}\cap\overline{\mathcal{D}}_{T}$, then for any $t\in[0,T]$, $(t,X_{t}^{*},Z_{t})\in{\cal O}_{T}^{c}\cap\overline{\mathcal{D}}_{T}$, $\mathbb{P}$-a.s.. Whether the capital injection is needed completely depends on the initial wealth level of the fund manager and initial level of the benchmark process.
@@
In what follows, we use the notation ${\cal O}_{T}^{f}$ to highlight the dependence of the capital injection region ${\cal O}_{T}$ defined by (3.12) on the function $f\in{\tt{C}}([0,T])$. Based on Corollary 4.4, for any $(x,z)\in\mathds{R}_{+}^{2}$ , let us consider the following mapping ${\cal I}:{\tt{C}}([0,T])\mapsto{\tt{C}}([0,T])$ given by
@@
A main challenge of this problem lies in the piecewise structure of the mapping (4.8), whose expression and the boundary both dependent on the input $f$. The free boundary that separates the capital injection region from the non-injection region is not fixed but evolves with the input $f$, creating a moving boundary problem in the fixed point analysis.
@@
Let $(x,z)\in\mathds{R}_{+}^{2}$. Then, there exists a fixed point $f^{*}\in{\tt{C}}([0,T])$ to the mapping (4.8) satisfying $f^{*}(t)>0$ for all $t\in[0,T]$. Let $\theta^{*,f^{*}}=(\theta_{t}^{*,f^{*}})_{t\in[0,T]}$ be defined in Theorem 3.4 for the fixed point $f^{*}\in{\tt{C}}([0,T])$, and define that, for $t\in[0,T]$,
$A_{t}^{*,f^{*}}:=0\vee\sup_{s\leq t}\left(\lambda\int_{0}^{s}f^{*}(u)du+(1-\lambda)Z_{s}-V_{s}^{\theta^{*,f^{*}}}\right),$
where $V^{\theta^{*,f^{*}}}=(V_{t}^{\theta^{*,f^{*}}})_{t\in[0,T]}$ satisfies SDE (2.6) under the strategy $\theta^{*,f^{*}}$ and $Z=(Z_{t})_{t\in[0,T]}$ is given by (2.7). Then, $(f^{*},\theta^{*,f^{*}},A^{*,f^{*}})=(f^{*}(t),\theta_{t}^{*,f^{*}},A_{t}^{*,f^{*}})_{t\in[0,T]}$ is a MFE.
@@
Obviously, $f^{*}\in{\tt{C}}([0,T])$ and $f^{*}(\cdot)>0$. As a consequence, we can easily verify that the function $f^{*}$ given by (4.15) is a fixed point to the mapping (4.8). This yields that $f^{*}$ given by (4.15) satisfies the consistency condition (4.1), and hence $(f^{*},\theta^{*,f^{*}},A^{*,f^{*}})$ is a MFE. Thus, the proof of the theorem is completed. ∎
@@ 5 Discussions on MFE
We present in this section some quantitative properties and numerical results of the MFE obtained in Theorem 4.8. We first recall that, for $(z,T)\in\mathds{R}_{+}\times(0,+\infty)$, the capital injection threshold of initial wealth is given by
@@
$\theta^{*,f^{*}}(t,x,z)$ $=\frac{(1-\lambda)\sigma_{Z}\mu}{\sigma}e^{\eta(T-t)}z,$ (5.2)
and the capital injection strategy $A^{*,f^{*}}_{t}=0$ a.s. for $t\in[0,T]$. In this case, the representative agent does not need to inject any capital.
We next discuss two extreme cases. As the competition weight parameter $\lambda$ tends towards $1$, the constraint condition in (2.4) becomes $\overline{V}_{t}^{n}\leq V_{t}^{i}+A_{t}^{i}$. This indicates that the agent is primarily concerned with their relative performance compared to other peers, resulting in the threshold $\hat{x}_{0}(z)\equiv 0$ and optimal portfolio $\theta^{*,f^{*}}(t,x,z)\equiv 0$. Without the pressure from the market index performance, the agent refrains from making any investments. Conversely, as the competition weight parameter $\lambda$ tends towards 0, the constraint condition in (2.4) becomes $Z_{t}^{i}\leq V_{t}^{i}+A_{t}^{i}$. In this scenario, the agent’s goal is to outperform the market index benchmark $Z_{t}^{i}$, disregarding the wealth performance of other agents. Consequently, the MFG problem reduces to the stochastic control problem discussed in Section 6 of Bo et al. (2021) with $\hat{x}_{0}(z)=(e^{\eta T}-1)z$ and $\theta^{*,f^{*}}(t,x,z)=\sigma_{Z}\mu/\sigma e^{\eta(T-t)}z$.
We next consider the sensitivity analysis on the return rate parameter $\mu_{Z}$ of the market index process. From (5.1), (5.2), and the relationship $\eta=\mu_{Z}-\mu\sigma_{Z}/\sigma$, it follows that $\partial\hat{x}_{0}/\partial\mu_{Z}>0$, and $\partial\theta^{*,f^{*}}/\partial\mu_{Z}>0$, which implies that both the threshold and the portfolio feedback function increase with respect to $\mu_{Z}$. When the market index benchmark process has a higher return, the agent will require a higher level of initial wealth and will invest more in the risky asset to outperform the targeted benchmark and avoid capital injection.
If the initial wealth level $x<\hat{x}_{0}(z)$, the capital injection is necessary. In this case, we can numerically calculate the MFE and value function. Our numerical method proceeds through the following key steps by leveraging the duality established in our theoretical analysis:
@@
    Initial capital injection assessment: we compute the capital injection threshold $\hat{x}_{0}(z)$ via (5.1). If $x_{0}\geq\hat{x}_{0}(z)$, the agent starts in the non-capital injection region. In this case, the optimal portfolio feedback function $\theta^{*,f^{*}}$ is computed directly via (5.2), no capital injection occurs ($A^{*,f^{*}}\equiv 0$), and the value function satisfies $v(t,x,z)=0$. Otherwise, we proceed to the following steps.
@@
Then, we plot the portfolio feedback function $x\mapsto\theta^{*,f^{*}}(t,x,z)$ in Figure 4 and value function (i.e., the expectation of the total optimal discounted capital injection) $x\mapsto v(t,x,z)$ in Figure 5. In particular, it is interesting to see that $\theta^{*,f^{*}}(t,x,z)$ is a decreasing function of $x$ for the fixed $t$ and $z$, which seems counter-intuitive. However, we note that the large value of $x$ indicates that the current wealth level of all agents or the population is also high, i.e., $f^{*}(t)$ is very large as well. As a result, in the equilibrium state, it has a larger chance for the agent to inject capital more frequently due to the tracking constraint. To reduce the cost of capital injection, the representative agent in the equilibrium will strategically reduce the allocation in the risky asset when $x$ becomes large to pull down $f^{*}(t)$ to lower the growth rate of the benchmark process. From Figure 3, it can be observed that under different initial values of the state process $X$, although $f^{*}(t)$ is always non-negative, its monotonicity in terms of $t$ may vary significantly under the impact of initial wealth. In particular, in the situation when $f^{*}(t)$ climbs to a very large level at some time $t$, we can see from Figure 3 that $f^{*}(t)$ will start to decrease afterwards, which echos from the observation in Figure 4 that the portfolio strategy under MFE is decreasing in $x$ such that the resulting $f^{*}(t)$ under MFE will decline when it hits a high level, that is, the representative agent will choose the portfolio strategy to maintain the growth rate of the wealth around a reasonable level.
@@
We next conduct some numerical examples on the sensitivity analysis with respect to some model parameters. Recall that the value function is an equivalent characterization of the expectation of the total optimal discounted capital injection. We illustrate in Figures 6 and Figure 7 the sensitivity of the expectation of the total optimal discounted capital injection as well as the optimal portfolio feedback function with respect to the competition weight parameter $\lambda$. All other model parameters are kept as in the baseline specification before. Consistent with the case without capital injections, the portfolio feedback function is decreasing in $\lambda$. As the competition weight increases, the agent places greater emphasis on relative performance against peers and correspondingly less on outperforming the market index in absolute terms. This shift in objective leads to a more conservative optimal allocation to the risky asset: the manager reduces exposure to risk and relies less on aggressive trading to generate outperformance. In parallel, Figure 6 shows that a smaller competition weight induces the manager to inject more capital to surpass the targeted benchmark; as $\lambda$ rises, the expected discounted injections decline. Relaxed competition concerns against peers (lower $\lambda$) encourage risk-taking that may boost short-term outperformance but raise the probability of drawdowns requiring injections. These comparative statics suggest that tuning $\lambda$ can serve as a policy lever. Institutions seeking capital discipline and lower risk budgets should adopt higher competition weights, whereas those prioritizing the performance of market index may allow lower $\lambda$, accepting higher expected capital injection.
@@
We also examine how the expectation of the total optimal discounted capital injection and the optimal portfolio feedback respond to changes in the volatility parameter $\sigma_{Z}$ of the market index process. Figures 8 and 9 plot these sensitivity results. The most prominent finding, illustrated in Figure 8, is that higher market index volatility $\sigma_{Z}$ leads to reduced expected capital injection. Recall that the effective benchmark is a linear combination $\lambda\overline{V}_{t}+(1-\lambda)Z_{t}$. When $\sigma_{Z}$ increases, the exogenous component $Z_{t}$ becomes more volatile, introducing greater variability into the benchmark itself. Faced with a more volatile benchmark, the agent’s optimal strategy naturally adapts to maintain a more conservative position relative to the benchmark. Mathematically, this translates to a decrease in the activity of the reflection process $L_{t}$, which represents cumulative capital injection. The manager’s adjusted strategy results in fewer instances where the portfolio value approaches the critical threshold requiring intervention.
@@
Figure 9 further reveals the non-monotonic property of the optimal portfolio feedback function with respect to $\sigma_{Z}$. Specifically, at high wealth levels, an increase in the market index volatility $\sigma_{Z}$ can lead to more aggressive portfolio weights, as the manager seeks to exploit higher volatility for potential outperformance, thereby tracking the market index and the strong performance of peers. Conversely, at low wealth levels, the same increase in $\sigma_{Z}$ prompts a reduction in portfolio risk to avoid triggering costly capital injections. In this scenario, the agent reduces investment to mitigate risk and may resort to direct capital injection to meet the benchmark requirement. From a practical perspective, these findings suggest that agent operating in more volatile market environments may benefit from lower expected capital injection costs. However, this comes at the potential cost of reduced opportunity to significantly outperform the benchmark during market upswings. The optimal strategy thus represents a careful balancing act between protection and performance that adapts rationally to the market condition.
Figure 6: The expectation of the total optimal discounted capital injection. The parameters are set to be $(x_{0},z_{0})=(2.02,20)$, $(t,z)=(0.5,20)$.
@@
Figure 8: The expectation of the total optimal discounted capital injection. The parameters are set to be $(x_{0},z_{0})=(2.02,20)$, $(t,z)=(0.5,20)$.

### v3

@@ 3 Equivalent Problem Formulation for Best Response Control
To tackle the best response control of the representative agent in problem (2.5), we introduce a new controlled state process to replace the original state process. Let $K_{t}:=\lambda\overline{V}_{t}+(1-\lambda)Z_{t}-V_{t}$ and $L_{t}:=0\vee\sup_{s\leq t}K_{s}$ for $t\in[0,T]$. Then, we define a new state process with reflection that $X_{t}:=L_{t}-K_{t}$, and hence
@@
$u(t,x,z):=\sup_{\theta\in\mathbb{U}_{t}}\mathbb{E}\left[-\int_{t}^{T}e^{-\rho(s-t)}dL_{s}^{X}\big|X_{t}=x,Z_{t}=z\right],$ (3.2)
where the admissible control set $\mathbb{U}_{t}$ is specified as the set of $\mathbb{F}$-adapted real-valued control processes $\theta=(\theta_{s})_{s\in[t,T]}$ satisfying $\mathbb{E}[\int_{t}^{T}|\theta_{s}|^{2}ds]<\infty$. Obviously, we have $\mathbb{U}=\mathbb{U}_{0}$ and the following equivalence between problems (2.5) and (3.2):
@@
We refer to $\mathcal{O}_{T}$ as the underperforming region and $\mathcal{O}_{T}^{c}\cap\overline{\mathcal{D}}_{T}$ as the outperforming region. In the underperforming region, even under the optimal portfolio strategy, the agent’s wealth falls below the benchmark process. In contrast, in the outperforming region, the agent can adopt a portfolio strategy such that the resulting wealth process can always outperform the benchmark. Moreover, Theorem 3.3 yields that the next corollary on the semi-explicit form of the optimal portfolio (feedback) function given by
@@ 4 Mean Field Equilibrium
and establishes the existence of a positive function $f^{*}\in{\tt{C}}([0,T])$ satisfying (4.1). Consequently, the MFE is given by $(f^{*},\theta^{*,f^{*}})=(f^{*}(t),\theta_{t}^{*,f^{*}})_{t\in[0,T]}$ characterized in Definition 2.1. On the other hand, it follows from (2.6) that $V_{t}^{\theta^{*,f}}=\mathrm{v}+\int_{0}^{t}\theta_{s}^{*,f}\mu ds+\int_{0}^{t}\theta_{s}^{*,f}\sigma dW_{s}$ for $t\in[0,T]$. Then, taking the expectation on both sides of this equation, we have
@@
Lemma 4.2 shows that the underperforming region $\cal O$} and outperforming region ${\cal O}_{T}^{c}\cap\overline{\mathcal{D}}_{T}$ are not interconnected. In other words, if the initial state $(0,x,z)\in{\cal O}_{T}$, $(t,X_{t}^{*},Z_{t})\in{\cal O}_{T}$, $\mathbb{P}$-a.s., for any $t\in[0,T]$; if $(0,x,z)\in{\cal O}_{T}^{c}\cap\overline{\mathcal{D}}_{T}$, then for any $t\in[0,T]$, $(t,X_{t}^{*},Z_{t})\in{\cal O}_{T}^{c}\cap\overline{\mathcal{D}}_{T}$, $\mathbb{P}$-a.s.. Whether the wealth process falls below the benchmark completely depends on the initial wealth level of the fund manager and initial level of the benchmark process.
@@
In what follows, we use the notation ${\cal O}_{T}^{f}$ to highlight the dependence of the underperforming region ${\cal O}_{T}$ defined by (3.11) on the function $f\in{\tt{C}}([0,T])$. Based on Corollary 4.4, for any $(x,z)\in\mathds{R}_{+}^{2}$ , let us consider the following mapping ${\cal I}:{\tt{C}}([0,T])\mapsto{\tt{C}}([0,T])$ given by
@@
A main challenge of this problem lies in the piecewise structure of the mapping (4.8), whose expression and the boundary both dependent on the input $f$. The free boundary that separates the underperforming region from the outperforming region is not fixed but evolves with the input $f$, creating a moving boundary problem in the fixed point analysis.
@@
Let $(x,z)\in\mathds{R}_{+}^{2}$. Then, there exists a fixed point $f^{*}\in{\tt{C}}([0,T])$ to the mapping (4.8) satisfying $f^{*}(t)>0$ for all $t\in[0,T]$. Let $\theta^{*,f^{*}}=(\theta_{t}^{*,f^{*}})_{t\in[0,T]}$ be defined in Theorem 3.3 for the fixed point $f^{*}\in{\tt{C}}([0,T])$. Then, $(f^{*},\theta^{*,f^{*}})=(f^{*}(t),\theta_{t}^{*,f^{*}})_{t\in[0,T]}$ is a MFE.
@@
Obviously, $f^{*}\in{\tt{C}}([0,T])$ and $f^{*}(\cdot)>0$. As a consequence, we can easily verify that the function $f^{*}$ given by (4.15) is a fixed point to the mapping (4.8). This yields that $f^{*}$ given by (4.15) satisfies the consistency condition (4.1), and hence $(f^{*},\theta^{*,f^{*}})$ is a MFE. Thus, the proof of the theorem is completed. ∎
@@ 5 Discussions on MFE
We present in this section some quantitative properties and numerical results of the MFE obtained in Theorem 4.8. We first recall that, for $(z,T)\in\mathds{R}_{+}\times(0,+\infty)$, the outperforming threshold of initial wealth is given by
@@
$\theta^{*,f^{*}}(t,x,z)$ $=\frac{(1-\lambda)\sigma_{Z}\mu}{\sigma}e^{\eta(T-t)}z.$ (5.2)
In this case, the representative agent’s wealth process under optimal portfolio strategy consistently outperforms the benchmark.
We next discuss two extreme cases. As the competition weight parameter $\lambda$ tends towards $1$, the benchmark process for agent $i$ becomes $\overline{V}_{t}^{n}$. This indicates that the agent is primarily concerned with their relative performance compared to other peers, resulting in the threshold $\hat{x}_{0}(z)\equiv 0$ and optimal portfolio $\theta^{*,f^{*}}(t,x,z)\equiv 0$. Without the pressure from the market index performance, the agent refrains from making any investments. Conversely, as the competition weight parameter $\lambda$ tends towards 0, the benchmark process for agent $i$ becomes $Z_{t}^{i}$. In this scenario, the agent’s goal is to outperform the market index benchmark $Z_{t}^{i}$, disregarding the wealth performance of other agents. Consequently, the MFG problem reduces to the stochastic control problem discussed in Section 6 of Bo et al. (2021) with $\hat{x}_{0}(z)=(e^{\eta T}-1)z$ and $\theta^{*,f^{*}}(t,x,z)=\sigma_{Z}\mu/\sigma e^{\eta(T-t)}z$.
We next consider the sensitivity analysis on the return rate parameter $\mu_{Z}$ of the market index process. From (5.1), (5.2), and the relationship $\eta=\mu_{Z}-\mu\sigma_{Z}/\sigma$, it follows that $\partial\hat{x}_{0}/\partial\mu_{Z}>0$, and $\partial\theta^{*,f^{*}}/\partial\mu_{Z}>0$, which implies that both the threshold and the portfolio feedback function increase with respect to $\mu_{Z}$. When the market index benchmark process has a higher return, the agent will require a higher level of initial wealth and will invest more in the risky asset to outperform the targeted benchmark.
If the initial wealth level $x<\hat{x}_{0}(z)$, even under the optimal portfolio strategy, the agent’s wealth falls below the benchmark process. In this case, we can numerically calculate the MFE and value function. Our numerical method proceeds through the following key steps by leveraging the duality established in our theoretical analysis:
@@
    Initial outperforming assessment: we compute the outperforming threshold $\hat{x}_{0}(z)$ via (5.1). If $x_{0}\geq\hat{x}_{0}(z)$, the agent starts in the outperforming region. In this case, the optimal portfolio feedback function $\theta^{*,f^{*}}$ is computed directly via (5.2), and the value function satisfies $v(t,x,z)=0$. Otherwise, we proceed to the following steps.
@@
Then, we plot the portfolio feedback function $x\mapsto\theta^{*,f^{*}}(t,x,z)$ in Figure 4 and value function $x\mapsto v(t,x,z)$ in Figure 5. In particular, it is interesting to see that $\theta^{*,f^{*}}(t,x,z)$ is a decreasing function of $x$ for the fixed $t$ and $z$, which seems counter-intuitive. However, we note that the large value of $x$ indicates that the current wealth level of all agents or the population is also high, i.e., $f^{*}(t)$ is very large as well. As a result, in the equilibrium state, it has a larger chance for the agent to falls below the benchmark process. more frequently due to the tracking constraint. To reduce the largest shortfall with reference to the benchmark process, the representative agent in the equilibrium will strategically reduce the allocation in the risky asset when $x$ becomes large to pull down $f^{*}(t)$ to lower the growth rate of the benchmark process. From Figure 3, it can be observed that under different initial values of the state process $X$, although $f^{*}(t)$ is always non-negative, its monotonicity in terms of $t$ may vary significantly under the impact of initial wealth. In particular, in the situation when $f^{*}(t)$ climbs to a very large level at some time $t$, we can see from Figure 3 that $f^{*}(t)$ will start to decrease afterwards, which echos from the observation in Figure 4 that the portfolio strategy under MFE is decreasing in $x$ such that the resulting $f^{*}(t)$ under MFE will decline when it hits a high level, that is, the representative agent will choose the portfolio strategy to maintain the growth rate of the wealth around a reasonable level.
@@
We next conduct some numerical examples on the sensitivity analysis with respect to some model parameters. Recall that the value function is an equivalent characterization of the expectation of the largest shortfall. We illustrate in Figures 6 and Figure 7 the sensitivity of the expectation of thelargest shortfall as well as the optimal portfolio feedback function with respect to the competition weight parameter $\lambda$. All other model parameters are kept as in the baseline specification before. Consistent with the outperforming case, the portfolio feedback function is decreasing in $\lambda$. As the competition weight increases, the agent places greater emphasis on relative performance against peers and correspondingly less on outperforming the market index in absolute terms. This shift in objective leads to a more conservative optimal allocation to the risky asset: the manager reduces exposure to risk and relies less on aggressive trading to generate outperformance. In parallel, Figure 6 shows that a smaller competition weight induces the manager to incur a larger expected largest shortfall relative to the benchmark;; as $\lambda$ rises, the expected largest shortfall decline. Relaxed competition concerns against peers (lower $\lambda$) encourage risk-taking that may boost short-term outperformance but raise the probability of drawdowns, which are reflected in a larger expected largest shortfall.. These comparative statics suggest that tuning $\lambda$ can serve as a policy lever. Institutions seeking lower risk budgets should adopt higher competition weights, whereas those prioritizing the performance of market index may allow lower $\lambda$, accepting a higher expected largest shortfall.
@@
We also examine how the expectation of the largest shortfall and the optimal portfolio feedback respond to changes in the volatility parameter $\sigma_{Z}$ of the market index process. Figures 8 and 9 plot these sensitivity results. The most prominent finding, illustrated in Figure 8, is that higher market index volatility $\sigma_{Z}$ leads to reduced expected largest shortfall. Recall that the effective benchmark is a linear combination $\lambda\overline{V}_{t}+(1-\lambda)Z_{t}$. When $\sigma_{Z}$ increases, the component $Z_{t}$ becomes more volatile, introducing greater variability into the benchmark itself. Faced with a more volatile benchmark, the agent’s optimal strategy naturally adapts to maintain a more conservative position relative to the benchmark. Mathematically, this translates to a decrease in the activity of the reflection process $L_{t}$, which represents largest shortfall. The manager’s adjusted strategy results in fewer instances where the portfolio value approaches the critical threshold requiring intervention.
@@
Figure 9 further reveals the non-monotonic property of the optimal portfolio feedback function with respect to $\sigma_{Z}$. Specifically, at high wealth levels, an increase in the market index volatility $\sigma_{Z}$ can lead to more aggressive portfolio weights, as the manager seeks to exploit higher volatility for potential outperformance, thereby tracking the market index and the strong performance of peers. Conversely, at low wealth levels, the same increase in $\sigma_{Z}$ prompts a reduction in portfolio risk to avoid incurring a costly largest shortfall. In this scenario, the agent reduces investment to mitigate risk and may accept a large shortfall to meet the benchmark requirement. From a practical perspective, these findings suggest that agent operating in more volatile market environments may benefit from a lower expected largest shortfall. However, this comes at the potential cost of reduced opportunity to significantly outperform the benchmark during market upswings. The optimal strategy thus represents a careful balancing act between protection and performance that adapts rationally to the market condition.
Figure 6: The expectation of the largest shortfall. The parameters are set to be $(x_{0},z_{0})=(2.02,20)$, $(t,z)=(0.5,20)$.
@@
Figure 8: The expectation of the largest shortfall. The parameters are set to be $(x_{0},z_{0})=(2.02,20)$, $(t,z)=(0.5,20)$.